If P = 50 watt and R = 2 ohms, then I = _
1
Expert's answer
2020-05-18T09:59:47-0400

Electric power is equal to product of electric current I and voltage V

"P=I\cdot V."

We know that according to Ohm's law "V=I\cdot R," so "P=I^2\cdot R."

If R=2 ohms and P = 50 watt, we get

"I = \sqrt{\dfrac{P}{I}} = \sqrt{\dfrac{50}{2}} = 5 \,\mathrm{A}."